1.verify that f(x)=2x/k(k+1) for x=0,1,2,.....k. can serve as a pmf of a random variable x

2.A fair coin is flipped until a head appears. Let N represent the number of tosses required to realise a head. Find the pmf of N

3.The pmf of a discrete random variable X is given by p(X=x)=kx for x=1,2,3,4,5,6.

Find the value of the constant k, p(X<4) and p(3<x<6)

4.A die is loaded such the probability that of a face showing up is proportional to the face number. Determine the probability of each sample point

5.Roll a fair die and let X be the square of the show up.Write the probability distribution of x hence compute p(X<15) and p(3<x<30)

6.Let x be a random variable the number of hours observed when two dice are rolled together once. show that x is a discrete random variable.

7.For each of the following determine c so that the function can serve as a pmf of a random variable X

  1. F(X)=cx for x=1,2,3,4,5
  2. f(x)=cx2 for x=0,1,2,...k
  3. f(x)=c(1/6)x for x=0,1,2,3,...
  4. f(x)=c2-x for x=0,1,2,....
1
Expert's answer
2021-07-16T16:28:46-0400

1.


"f(x)=x(\dfrac{2}{k(k+1)})\geq 0, x=0,1,2,..."


"\displaystyle\sum_{x=0}^kf(x)=\displaystyle\sum_{x=0}^kx(\dfrac{2}{k(k+1)})"

"=\dfrac{2}{k(k+1)}(\dfrac{k(k+1)}{2})=1"

"f(x)=x(\dfrac{2}{k(k+1)})\geq 0, x=0,1,2,..." can serve as a pmf of a random variable "x."


2.

A geometric random variable X with parameter p has probability mass function

"f(N) = p(1\u2212 p) ^N, N = 0,1,2,... ."

Given "p=0.5"


"f(N) = 0.5(1\u2212 0.5) ^N, N = 0,1,2,... ."

"f(N) = 0.5 ^{N+1}, N = 0,1,2,... ."

3.


"k(1)+k(2)+k(3)+k(4)+k(5)+k(6)=1"

"k=\dfrac{1}{21}"

"P(X<4)=\dfrac{1}{21}(1+2+3)=\dfrac{2}{7}"

"P(3<X<6)=\dfrac{1}{21}(4+5)=\dfrac{3}{7}"

4.

"k+2k+3k+4k+5k+6k=1"

"k=\dfrac{1}{21}"

"P(X=1)=\dfrac{1}{21}"

"P(X=2)=\dfrac{2}{21}"

"P(X=3)=\dfrac{1}{7}"

"P(X=4)=\dfrac{4}{21}"

"P(X=5)=\dfrac{5}{21}"

"P(X=6)=\dfrac{2}{7}"

5.


"\begin{matrix}
 x & 1 & 4 & 9 & 16 & 25 & 36 \\
 p(x) & 1/6 & 1/6 & 1/6 & 1/6 & 1/6 & 1/6
\end{matrix}"

"p(X<15)=\dfrac{1}{6}+\dfrac{1}{6}+\dfrac{1}{6}=\dfrac{1}{2}"


"p(3<X<30)=\dfrac{1}{6}+\dfrac{1}{6}+\dfrac{1}{6}+\dfrac{1}{6}=\dfrac{2}{3}"

6.A discrete random variable is an rv whose possible values either constitute a

finite set or else can be listed in an infinite sequence in which there is a first

element, a second element, and so on (“countably” infinite).

A random variable the number of hours observed when two dice are rolled together once is

 countable number of possible values.

Therefore this random variable is discrete.


7.


1.

"c+2c+3c+4c+5c=1"

"c=\dfrac{1}{15}"

2.

"\displaystyle\sum_{x=0}^{k}cx^2=\dfrac{ck(k+1)(2k+1)}{6}=1"

"c=\dfrac{1}{k(k+1)(2k+1)}"



3.

"\displaystyle\sum_{x=0}^{\infin}c(\dfrac{1}{6})^x=\dfrac{c}{1-\dfrac{1}{6}}=\dfrac{6c}{5}=1"

"c=\dfrac{5}{6}"

4.

"\displaystyle\sum_{x=0}^{\infin}c(2)^{-x}=\displaystyle\sum_{x=0}^{\infin}c(\dfrac{1}{2})^x=\dfrac{c}{1-\dfrac{1}{2}}=2c=1"

"c=\dfrac{1}{2}"